OpenStudy (anonymous):

Find the equation of the tangent to the curve y=x+sinx at the point where x=π/2

OpenStudy (anonymous):

Using calculus you get z=ax+b, with a = derivative of y, wich is 1+cos(x), at the point x=pi/2 you have 1 because cos(pi/2)=0, substituting x in the equation of y, you get y=pi/2+1, so substituting in z we get pi/2+1=pi/2+b, so b = 1, and the tangent is z=x+1

OpenStudy (anonymous):

Thank you so much!!! :D